OpenStudy (idealist10):

Find the radius of curvature for y=e^x at x=1.

OpenStudy (idealist10):

\[k(x)=\frac{ \left| y'' \right| }{ [1+(y')^2]^{3/2} }\]

OpenStudy (idealist10):

\[k(x)=\frac{ e^x }{ (e ^{2x}+1)^{3/2} }\]

OpenStudy (idealist10):

^ That above is the curvature, so how do I find the radius of curvature at x=1?

myininaya (myininaya):

plug in 1

OpenStudy (irishboy123):

normally, radius of curvature is 1/k

OpenStudy (idealist10):

So I got \[k(1)=\frac{ e }{ (e^2+1)^{3/2} }\]

myininaya (myininaya):

http://mathworld.wolfram.com/RadiusofCurvature.html this site says it should be the reciprocal of that

myininaya (myininaya):

which I think that @IrishBoy123 said what he said above

myininaya (myininaya):

why I think*

OpenStudy (idealist10):

So I got \[R=\frac{ (e^2+1)^{3/2} }{ e }\]

myininaya (myininaya):

that sounds great to me :)

OpenStudy (idealist10):

Thank you so much, @myininaya !